raul has a younger sister who is four years younger the sum combined age is 28
Answer by ankor@dixie-net.com(22740)

l has a younger sister who is four years younger
r = s + 4
r - s = 4
the sum combined age is 28.
r + s = 28
use elimination on these two equations
r - s = 4
r + s = 28
---------------adding eliminates s, find r
2r = 32
r = 16 is R's age
I'll let you find the sister's age. Confirm that they add up to 28